Analyze the key characteristics of the following quadratic function. f(x)=x2-9 Based on the graph, which statement is not true? 0 A. The y-intercept is the same point as the vertex O B. The graph increases in the interval 0 to oo ° C. The graph is always positive. O D. The zeros of this function are at 3 and -3
Answer #1

f(x)=x²-9

when we put x =0, then

f(x) = 0² -9 = 0 -9 = -9 which is negative. So graph is NOT always positive.

Correct option is C

) 049 987654321